OpenStudy (anonymous):

F(s)=7s^2+23s+30/(s-2)(s^2+2s+5) Inverse laplace transform just wondering if i should start it like....

OpenStudy (anonymous):

complete the square for the \[s^2+2s+5\] in the denominator giving \[(s+1)^2+4\]

OpenStudy (anonymous):

but partial fraction decomposition would give me C/4 and don't know if that's acceptable...

OpenStudy (anonymous):

@abb0t assistance please???

OpenStudy (abb0t):

Wait, what?

OpenStudy (anonymous):

just wondering if C/4 would be acceptable for partial fraction decomp...

OpenStudy (abb0t):

\(\frac{c}{4}\) ?

OpenStudy (abb0t):

Oh, yeah. It is. It's like \(\frac{1}{4}C\) like multiplied by constant.

OpenStudy (anonymous):

oh ok cool.. thanks!

OpenStudy (anonymous):

@oldrin.bataku if you get a chance and could help me that would be awesome. Im not even sure if im starting this one right

OpenStudy (anonymous):

@abb0t could you help with this one? i don't think how i tried to start it is right...

OpenStudy (anonymous):

and if it is then i just messed up the partial fractions

OpenStudy (abb0t):

Hmm..ok. Well, you could use partial fraction decomp, and substitute (s+1)\(\rightarrow\)s correct?

OpenStudy (anonymous):

s+1?

OpenStudy (anonymous):

from completing the squre?

OpenStudy (anonymous):

oh do i have to take derivative of something?

OpenStudy (zzr0ck3r):

is latex showing up for you guys @abb0t @rperez36 ?

OpenStudy (abb0t):

I think so. Yeah.

OpenStudy (zzr0ck3r):

word ty

OpenStudy (abb0t):

You actually could separate each function. Yes.

OpenStudy (anonymous):

\[\frac{ A }{(s-2) }+\frac{ Bs+C }{ (s^2+2s+5) }\] so u find the constants A, B and C then take inverse laplace of each fraction

OpenStudy (abb0t):

You could use the derivative on the last one I believe. I forgot the theorems and corollaries..

OpenStudy (anonymous):

ah ok @litchlani that should be it. i'll try it and see if i can derive the answer.

OpenStudy (anonymous):

@litchlani i got close to the answer but not quite. Was definitely the way to approach it though!

OpenStudy (anonymous):

sorry I wasn't home! $$F(s)=\frac{7s^2+23s+30}{(s-2)(s^2+2s+5)}=\frac{A}{s-2}+\frac{Bs+C}{s^2+2s+5}\\A(s^2+2s+5)+(Bs+C)(s-2)=7s^2+23s+30\\(A+B)s^2+(2A-2B+C)s+(5A-2C)=7s^2+23s+30\\A+B=7\\2A-2B+C=23\\5A-2C=30$$ solving we use:$$A+B=7\\2A+2B=14\\2A-2B+C=23\\2A-2B+C+2A+2B=23+14\\4A+C=37\\8A+2C=74\\5A-2C=30\\13A=104\\A=8\\B=7-8=-1\\C=37-4(8)=5$$i.e. our decomposition yields:$$F(s)=\frac8{s-2}-\frac{s-5}{s^2+2s+5}$$note however that \(s^2+2s+5\) can be rewritten by completing the square \(s^2+2s+5=s^2+2s+1+4=(s+1)^2+2^2\). We can also rewrite \(s-5=s+1-6\):$$F(s)=\frac8{s-2}-\frac{s+1}{(s+1)^2+2^2}+3\frac2{(s+1)^2+2^2}$$

OpenStudy (anonymous):

now, the inverse Laplace is straightforward in this form:$$f(t)=\mathcal{L}^{-1}\left\{\frac8{s-2}-\frac{s+1}{(s+1)^2+2^2}+3\cdot\frac2{(s+1)^2+2^2}\right\}=8e^{2t}-e^{-t}\cos2t+3e^{-t}\sin2t$$

OpenStudy (anonymous):

\(f(t)=8e^{2t}-e^{-t}\cos2t+3e^{-t}\sin2t\)